General Informations

Status-Key

Status/green.gif

Resolved

Status/orange.gif

Still working but with some errors

Status/red.gif

Pending


MySQL Server downtime

STATUS: Status/green.gif

2013-11-18: 12:00
We are experiencing Problems with the MySQL database cluster. Affected are also all websites using the MySQL instance on remi.ee.ethz.ch
2013-11-18: 12:45
The system is working again.


Server oenone NFS problems

STATUS: Status/green.gif

2013-11-12: 04:00

We see NFS related problems on server oenone again. Affected were homes of all users from AUT, BIWI, IKT and VAW.

2013-11-12: 07:40
The system is again back online.


Server oenone NFS problems

STATUS: Status/green.gif

2013-10-16: 02:00

Starting from around 02:00 in the morning we had NFS related problems on server oenone. Affected were homes of all users from AUT, BIWI, IKT and VAW. Webpages of these institutes may also have been affected.

2013-10-16: 03:00
Server oenone recovered itself and is running as normal again.


Maintenance work on D-ITET's central IT infrastructure

STATUS: Status/green.gif

2013-09-28: 14:30
All services are back online
2013-09-28: 10:00-16:00
To keep our systems up to date with the newest software and security releases, we need to update our servers on a regular basis. For this reason we are going to install latest Oracle patches on our main servers. The servers will be rebooted during this maintenance

To prevent data corruption/loss please do the following:

More details:


Server oenone crash

STATUS: Status/green.gif

2013-09-04: 21:30

oenone nfs server crashed. Affecting users of AUT, BIWI, IKT and VAW. Homes and Webpages of these institutes.

2013-09-05: 04:00
we needed to reset the server to bring it again up.


Server oenone crash

STATUS: Status/green.gif

2013-06-15: 08:30

oenone nfs server crashed. Affecting users of AUT, BIWI, IKT and VAW. Homes and Webpages of these institutes.

2013-06-15: 11:45
we needed to reset the machine as the service for providing the home directories was unresponsive.


ETZSPEZ - HP 6100 encountered print quality issues

STATUS: Status/green.gif

2013-05-30 12:00
The Printer has been fixed.
2013-05-30 09:00
Technician is mending the printer right now. Printer is offline for the next 4 hours.
2013-05-16 12:00
An external technician has been informed and will fix the printer in the room ETZ J66 (ETZSPEZ) as soon as possible.


The Informatikdienste have some problems with their NAS

STATUS: Status/green.gif

2013-04-23 14:42
As a result of the disorders of the ID-NAS, the VPP printers don't work properly at the moment.
2013-04-24 07:00
All ID Services should be back to normal.


Outage Network Infrastructure ETH

STATUS: Status/green.gif

2013-03-06 11:55
Again Informatikdienste has a global problem with the network infrastructure.
2013-03-06 13:00

Services should be all back. We got informed about the cause by the Informatikdienste. A Hardware-Loadbalancer in RZ crashed and was in an undefined state. Therefore the failover to HCI did not work.


Maintenance work on D-ITET's central IT infrastructure

STATUS: Status/green.gif

2013-02-23: 16:00
All systems are back online.
2013-02-23: 10:00-14:00
To keep our systems up to date with the newest software and security releases, we need to update our servers on a regular basis. For this reason we are going to install latest Oracle patches on our main servers. The servers will be rebooted during this maintenance

To prevent data corruption/loss please do the following:

More details:


Outage Network Infrastructure ETH

STATUS: Status/green.gif

2013-02-15 08:05
Again Informatikdienste has a global problem with the network infrastructure.
2013-02-15 09:10
Network is coming back, still problems present.
2013-02-15 09:20
Network ist still not 100% recovered.
2013-02-15 09:45
Network completly down again. We still don't have an update from Informatikdienste what is going on.
2013-02-15 10:45
Network ist still not stable.
2013-02-15 12:10
Network is coming back to normal, we are working on restoring the services.
2013-02-15 12:30
Most of our services are back online.
2013-02-15 14:00

We got informed about the cause by the Informatikdienste. Basically a virtuall firewall which was not needed anymore was deleted by the Networkteam at ID. In consequence all interfaces of all virtual firrewalls got down bringing down a big part of the ETH network. Neither a reload of the firewall hardware helped on their side, so they needed to reinstall all appliances. See the status website for a german written explanation.


DNS Outage Network ETH

STATUS: Status/green.gif

2013-02-14 14:35
Currently DNS at ETH is down. ID will keep us informed. This affects all services.
2013-02-14 14:45
Services are stabilizing.


DNS Outage Network ETH

STATUS: Status/green.gif

2013-02-12 16:30
Currently DNS at ETH is down. ID will keep us informed. This affects all services.
2013-02-12 18:50
Services are stabilizing.
2013-02-12 19:30

Not all DNS names are yet restored. However most services affecting D-ITET customers should work again. More informations will be publisched at Informatikdienste Statusseite. In case you experience some specific problems please contact us at support@ee.ethz.ch


Server oenone crash

STATUS: Status/green.gif

2013-02-06: 23:00

oenone nfs server crashed. Affecting users of AUT, BIWI, IKT and VAW. Homes and Webpages of these institutes.

2013-02-07: 07:10
we needed to reset the machine as the service for providing the home directories was unresonsive.


Hanging people.ee.ethz.ch Webserver

STATUS: Status/green.gif

2013-02-05: 06:00
The webserver for people.ee.ethz.ch was hanging this morning and was not pingable anymore.
2013-02-05: 07:00
We reseted the webserver and now investigating the issue. It's probably kernel- and hardware related.


Server oenone crash

STATUS: Status/green.gif

2013-01-14: 21:28

oenone crashed. Affecting users of AUT, BIWI, IKT and VAW.

2013-01-14 21:52
oenone rebooted and is running again.


CategoryEDUC

Status/Archive/2013 (last edited 2016-01-06 13:07:06 by alders)